OpenStudy (anonymous):

Can anyone please help me. Points P(1, 5) and Q(4, 1) on a coordinate grid represent side PQ of triangle PQR. What is the length of side PQ of the triangle? 5 units 25 units 7 units 13 units

OpenStudy (anonymous):

Would the answer be 7 units?

OpenStudy (anonymous):

length of PQ : Distance formula : \[\sqrt{(x2-x1)^2 + (y2-y1)^2}\] \[\sqrt{(4-1)^2 +( 1-5)^2}\] \[\sqrt{25}\] \[= 25\]

OpenStudy (anonymous):

= 5 *

OpenStudy (anonymous):

Oh ok, thank you I see where I messed up. Thank you so much for explaining it.

OpenStudy (anonymous):

yep i bet you mess up with the distance formula... anyway hope i was helpful good luck

OpenStudy (anonymous):

You were very helpful thanks again.